OpenStudy (anonymous):

7n+6n-5=4n+4 solve for n?

OpenStudy (anonymous):

7n + 6n - 5 = 4n + 4 First start by adding up like terms; 7n + 6n = 13n Now we have 13n - 5 = 4n + 4 Now lets get the 4n on the right side of the equation on the left by subtracting 4n from both sides of the equation; we get: 9n - 5 = 4 Now lets get the constant out of the left side of the equation by adding 5 to both sides of the equation; 9n = 9 Now we have to get rid of the 9 in front of the n variable and we can do that by dividing both sides of the equation by 9, We get, n = 1 as our answer.

OpenStudy (anonymous):

13n-4n=4+5 9n=9 n=1

OpenStudy (anonymous):

Thank You Very Much -Andrewj &&' Leaper

OpenStudy (anonymous):

A medal would be nice XD haha